OpenStudy (anonymous):

Solve the equation 2^(5x+1) = 8^(x-2)

OpenStudy (anonymous):

8=2^3 2^(5x+1)=2^3^(x-2) 2^(5x+1)=2^[3(x-2)] 5x+1=3(x-2)

OpenStudy (anonymous):

if you need to show intermediate step between 2^(5x+1)=2^[3(x-2)], take log_2 of each side \[\log_{2}2^{5x+1}=\log_{2}2^{3(x-2)}\] \[(5x+1)\log_{2}2=[3(x-2)]\log_{2}2\] \[\log_{2}2=1\] \[5x+1=3(x-2)\]

OpenStudy (anonymous):

This was very helpful. Thank you very much!

OpenStudy (anonymous):

yw